# TextureSplat: Per-Primitive Texture Mapping for Reflective Gaussian Splatting

### v1

@@
Abstract.
Gaussian Splatting have demonstrated remarkable novel view synthesis performance at high rendering frame rates. Optimization-based inverse rendering within complex capture scenarios remains however a challenging problem. A particular case is modelling complex surface light interactions for highly reflective scenes, which results in intricate high frequency specular radiance components. We hypothesize that such challenging settings can benefit from increased representation power. We hence propose a method that tackles this issue through a geometrically and physically grounded Gaussian Splatting borne radiance field, where normals and material properties are spatially variable in the primitive’s local space. Using per-primitive texture maps for this purpose, we also propose to harness the GPU hardware to accelerate rendering at test time via unified material texture atlas.
1. Introduction
@@
Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) revolutionized the field by representing scenes as continuous neural implicit functions optimized through differentiable volume rendering. Building on this foundation, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) reintroduced point-based graphics by replacing neural networks with explicit 3D Gaussian primitives, achieving both real-time rendering and state-of-the-art quality. These primitives are rendered through volume resampling (Zwicker et al., 2002), with their parameters optimized via gradient descent-based differentiable rendering. More recently, 2D Gaussian Splatting (2DGS) (Huang et al., 2024) improved multi-view consistency by using planar Gaussian primitives that better align with surfaces.
Despite these advances, accurately representing highly reflective surfaces remains challenging. Reflective objects exhibit complex view-dependent effects that depend on surface normals, material properties, and environmental lighting. Recent methods such as RefGaussian (Yao et al., 2025) and 3DGS-DR (Ye et al., 2024) attempt to model these effects using per-Gaussian material properties and physically-based rendering approaches. However, they often struggle with high-frequency specular highlights and sharp reflections. We hypothesise that this is due in part to the inherent resolution limitations of using a single attribute value per primitive.
@@
Inspired by the distinction between Gouraud shading (Gouraud, 1971) (constant per-vertex attributes) and Phong shading (Phong, 1975) (interpolated attributes) in traditional computer graphics, we introduce per-primitive texture maps for material properties in 2D Gaussian splatting. This approach effectively decouples the geometric representation (Gaussian primitives) from the appearance representation (material textures), allowing us to model high-frequency material variations without increasing the number of primitives.
@@
2. Related Work
@@
Neural Radiance Fields (Mildenhall et al., 2020) (NeRFs) have been dominating the 3D shape and appearance modelling recently, based on the astounding success of implicit representations combined with differentiable volume rendering (Levoy, 1988; Max, 1995). They represent scenes using view-dependent radiance and density fields parameterized by MLPs. When density is modeled as a function of a signed distance field, NeRF variants enable more accurate geometry reconstruction (Wang et al., 2021; Yariv et al., 2021; Li et al., 2023; Wu et al., 2022). However, multi-scale volume rendering demands frequent MLP evaluations, limiting real-time performance. Grid-based methods (Müller et al., 2022; Sun et al., 2022; Fridovich-Keil et al., 2023, 2022; Chen et al., 2022; Sun et al., 2024) alleviate this but often struggle with large, unbounded scenes even with level-of-detail grids (Liu et al., 2020). Gaussian splatting (3DGS) (Kerbl et al., 2023) emerged lately as a strong alternative to NeRFs, offering state-of-the-art novel view synthesis and real time rendering frame rates. It extends the eliptical weighted average (EWA) volume resampling framework (Zwicker et al., 2002, 2001) to learnable (Kingma and Ba, 2015) inverse rendering, modelling scenes with explicit Gaussian kernel primitives, that can be sorted and rasterized efficiently. The 2DGS representation (Huang et al., 2024) leverages planar 2D primitives instead of volumetric ones (3DGS), and performs precise 2D kernel evaluation in object space as opposed to approximative ones in screen space (3DGS), thus leading to superior geometric modelling and multi-view consistency.
@@
Both vanilla NeRFs and 3DGS assume low-frequency view dependency. Hence, they can struggle with highly reflective scenes. One strategy to improve in this department is using shading functions that are reflection direction aware (Verbin et al., 2022; Ge et al., 2023; Jiang et al., 2024b). For instance, Ref-NeRF (Verbin et al., 2022) extends NeRFs with a new parameterization for view-dependent radiance and incorporates normal vector regularization. The shading function can be more physically grounded, and this enables additional application such as relighting and material editing. In this regard, other NeRF and GS based methods proposed to model light interaction using the explicit rendering equation with BRDF functions. NeRV (Srinivasan et al., 2021) models reflectance and visibility fields separately for accurate relighting and editing. NeRO (Liu et al., 2023) Learns surface normals and view-dependent radiance jointly using RGB and environment maps. EnviDR (Liang et al., 2023) Uses environment maps to capture spatially-varying reflections in neural rendering. The next wave of work tackled shading for Gaussian Splatting (e.g. (Tang and Cham, 2024; Zhu et al., 2024; Liang et al., 2024; Jiang et al., 2024a; Ye et al., 2024; Gao et al., 2024; Yao et al., 2025)). GShader (Jiang et al., 2024a) applies a simplified shading function on each fragment. 3DGS-DR (Ye et al., 2024) introduces deferred shading at pixel level, and stabilizes the optimization by smoothing out normal gradients. 3iGS (Tang and Cham, 2024) introduces tensor decomposition of spherical harmonics. GS-ROR (Zhu et al., 2024) uses signed distance function priors to improve specular rendering. (Gao et al., 2024) employes ray-tracing to model inter-reflections. It approximates the rendering equation with Monte Carlo sampling which reduces rendering speed. Building on the 2DGS representation, our baseline Ref-Gaussian (Yao et al., 2025) decomposes the scene into geometry, material and lighting through the split-sum approximation of the rendering equation while incorporating an indriect light attribute, enabling inter-reflections while being fast to render. It achieves the state-of-the-art performance on the standard reflective scene novel view synthesis benchmarks. We propose to improve of its physical material and normal representations.
@@
In parallel with our work, we noticed that a few papers (Xu et al., 2024; Chao et al., 2025; Huang and Gong, 2024) introduce the texture attribute representation for Gaussian Splatting as well. They use it to model a view independent component of the color. Differently, we explore this representation for rendering material properties and normals within 2DGS enabled physically based rendering to reconstruct challenging highly reflective scenes. Differently, we also propose leveraging texture atlases to enable hardware acceleration at test-time rendering.
3. Method
@@
In this section, we present our approach to enhance 2D Gaussian Splatting (2DGS) (Huang et al., 2024) for representing highly reflective 3D scenes. We first provide background on 2DGS, then introduce our per-primitive texture mapping method that enables high-frequency detail on flat Gaussian primitives. Finally, we describe our hardware-accelerated implementation using texture atlases and the physically-based rendering model we use for reflective scenes.
3.1. Background: 2D Gaussian Splatting
@@
We build upon 2D Gaussian Splatting (2DGS) (Huang et al., 2024), which represents a scene using oriented planar disks offering improved multi-view consistency compared to 3D Gaussian Splatting (Kerbl et al., 2023). Each primitive is characterized by its position , tangential vectors and , and scaling factors . The primitive’s normal is defined by the cross product .
@@
For a point in the local coordinate space, the 2D Gaussian value is evaluated as:
Given pixel , Gaussians are rendered through ray-primitive plane intersection computation (Zwicker et al., 2004):
where is the view transformation matrix. The intersection is computed efficiently by transforming two orthogonal planes defined by the ray into the splat’s local coordinate system:
This yields a closed-form solution for the intersection coordinates:
where denotes the -th component of the vector.
@@
3.2. Per-Primitive Texture Mapping
@@
The original Elliptical Weighted Average (EWA) splatting paper (Zwicker et al., 2002) assumed constant attributes within the projected kernel’s footprint for analytical screen-space filtering, a characteristic inherited by standard 3DGS (Kerbl et al., 2023). However, the 2DGS formulation we adopt (Huang et al., 2024) computes ray-splat intersections directly and does not rely on EWA-style screen-space filtering, making it naturally suited for spatially varying attributes defined on the splat itself.
@@
Memory efficiency: Fewer primitives means reduced memory usage, improved cache coherence, faster rasterization and sorting operations.
@@
3.3. Texture Mapping Implementation
@@
3.4. Hardware Acceleration via Texture Atlases
To efficiently render optimized scenes at test time with per-primitive textures, we leverage hardware-accelerated texture filtering by packing individual primitive textures into texture atlases. This approach is inspired by methods like Ptex (Burley and Lacewell, 2008) and seamless texture atlases (Purnomo et al., 2004), adapted to the specific needs of Gaussian splatting.
3.4.1. Texture Atlas Construction
Each splat is associated with a small, fixed-resolution texture (e.g., texels, where is typically or ). For splats and total channels for all textured attributes, this texel data is organized into one or more large 2D atlas textures. We pack the individual primitive textures into larger texture atlases that can be efficiently accessed by the GPU by determining the maximum atlas dimensions based on GPU capabilities, calculating how many textures fit along each dimension, computing the total number of textures per atlas, then determining the number of atlases needed.
3.4.2. Accelerated Texture Sampling
(1) For a given primitive and local coordinates , we map to normalized texture coordinates as described above.
(2) We retrieve the chart coordinates from the indirection buffer.
(3) We compute the atlas coordinates:
where and are the total number of charts along the width and height of the atlas, respectively.
(4) We sample the atlas using hardware-accelerated bilinear filtering with CUDA’s ‘tex2D()‘ intrinsic:
| Shiny Blender (Verbin et al., 2022) | Glossy Synthetic (Liu et al., 2023) | Real (Verbin et al., 2022) | ||||||||||||||||
@@ 3.4.2. Accelerated Texture Sampling
| LPIPS | Ours | 0.075 | 0.027 | 0.075 | 0.039 | 0.004 | 0.066 | 0.037 | 0.065 | 0.034 | 0.043 | 0.040 | 0.052 | 0.041 | 0.051 | 0.276 | 0.250 | 0.263 |
3.5. Physically-Based Deferred Rendering
Following the deferred rendering approach used in 3DGS-DR (Ye et al., 2024) and Ref-Gaussian (Yao et al., 2025), we first splat material attributes to screen-space buffers, then apply physically-based shading in a separate pass.
3.5.1. Material Properties
Each 2D Gaussian is associated with texture maps for the following material properties: Albedo , Metallic , Roughness , Tangent normal : Represents normal perturbations in the tangent space. For memory efficiency, we encode tangent normals using only two components and reconstruct the third component at runtime using .
3.5.2. Normal Mapping
3.5.3. Attribute Splatting
where represents the combined screen-space buffers:
3.5.4. Physically-Based Shading
The diffuse term writes:
The first term, , depends solely on the view angle and roughness, which we precompute and store in a 2D lookup texture. is the direct environment lighting queried from a learnable environment map in the reflection direction , using roughness for mipmap selection. is the blended indirect lighting component.
3.5.5. Inter-reflection Modeling
To handle inter-reflections, we separately model direct light (not blocked by scene elements) and indirect light (all other light). We approximate the visibility of incident light by ray-tracing an extracted mesh as in (Yao et al., 2025) to determine whether light is self-occluded along the reflected direction. For the indirect lighting component, each Gaussian is assigned an additional view-dependent color , modeled by spherical harmonics. During rendering, is evaluated in the reflected direction at the Gaussian level, and alpha blending is applied to aggregate the indirect lighting map.
3.6. Training
where is the RGB reconstruction loss with balancing weight . The normal consistency loss encourages alignment of the Gaussians with the surface by minimizing the cosine difference between the rendered normal and the surface normal derived from rendered depth. The edge-aware normal smoothness loss regularizes normal variation in regions of low texture, where is the ground truth image. During training, we periodically extract the object’s surface mesh using truncated signed distance function (TSDF) fusion to compute ray-surface intersections for visibility checks, accelerated using a bounding volume hierarchy (BVH). We use a single NVIDIA RTX A6000 GPU for our evaluations.
4. Experiments

### v2

@@
Abstract
Gaussian Splatting have demonstrated remarkable novel view synthesis performance at high rendering frame rates. Optimization-based inverse rendering within complex capture scenarios remains however a challenging problem. A particular case is modelling complex surface light interactions for highly reflective scenes, which results in intricate high frequency specular radiance components. We hypothesize that such challenging settings can benefit from increased representation power. We hence propose a method that tackles this issue through a geometrically and physically grounded Gaussian Splatting borne radiance field, where normals and material properties are spatially variable in the primitive’s local space. Using per-primitive texture maps for this purpose, we also propose to harness the GPU hardware to accelerate rendering at test time via unified material texture atlas. Code will be available at TextureSplat.
1 Introduction
@@
Neural Radiance Fields (NeRF) [nerf] revolutionized the field by representing scenes as continuous neural implicit functions optimized through differentiable volume rendering. Building on this foundation, 3D Gaussian Splatting (3DGS) [3dgs] reintroduced point-based graphics by replacing neural networks with explicit 3D Gaussian primitives, achieving both real-time rendering and state-of-the-art quality. These primitives are rendered through volume resampling [ewa], with their parameters optimized via gradient descent-based differentiable rendering. More recently, 2D Gaussian Splatting (2DGS) [2DGS] improved multi-view consistency by using planar Gaussian primitives that better align with surfaces.
Despite these advances, accurately representing highly reflective surfaces remains challenging. Reflective objects exhibit complex view-dependent effects that depend on surface normals, material properties, and environmental lighting. Recent methods such as RefGaussian [RefGaussian] and 3DGS-DR [3DGS-DR] attempt to model these effects using per-Gaussian material properties and physically-based rendering approaches. However, they often struggle with high-frequency specular highlights and sharp reflections. We hypothesize that this is due in part to the inherent resolution limitations of using a single attribute value per primitive.
@@
Inspired by the distinction between Gouraud shading [gouraud] (constant per-vertex attributes) and Phong shading [phong] (interpolated attributes) in traditional computer graphics, we introduce per-primitive texture maps for material properties in 2D Gaussian splatting. This approach effectively decouples the geometric representation (Gaussian primitives) from the appearance representation (material textures), allowing us to model high-frequency material variations without increasing the number of primitives.
@@
2 Related Work
2.1 Radiance Fields for 3D Scene Representations
@@
Neural Radiance Fields [nerf] (NeRFs) have been dominating the 3D shape and appearance modelling recently, based on the astounding success of implicit representations combined with differentiable volume rendering [levoy1988volume, volrend]. They represent scenes using view-dependent radiance and density fields parameterized by MLPs. When density is modeled as a function of a signed distance field, NeRF variants enable more accurate geometry reconstruction [neus, volsdf, neuralangelo, SparseCraft, GeoTransfer, voxurf]. However, multi-scale volume rendering demands frequent MLP evaluations, limiting real-time performance. Grid-based methods [ingp, dvgo, kplanes, plenoxels, tensorrf, sun2024sparse] alleviate this but often struggle with large, unbounded scenes even with level-of-detail grids [nsvf]. Implicit reconstruction has been made more robust to noise and sparse observations, whether from images or point clouds, through the use of generalizable data priors (e.g. [pixelnerf, mvsnerf, geonerf, genlf, fssdf, nksr, mixing, robust, convocc]) and a variety of regularization strategies (e.g. [regnerf, freenerf, dietnerf, dsnerf, rgbdnerf, dro, nap, sparseocc, robust, digs, igr]). Gaussian splatting (3DGS) [3dgs] emerged lately as a strong alternative to NeRFs, offering state-of-the-art novel view synthesis and real time rendering frame rates. It extends the elliptical weighted average (EWA) volume resampling framework [ewa, zwicker2001ewa] to inverse rendering, modelling scenes with explicit Gaussian kernel primitives, that can be sorted and rasterized efficiently. Recent extensions of Gaussian splatting include building generalizable models [pixsplat, mvsgaussian, SparSplat, HiSplat], bundle-adjustment–based formulations [bad, colfree, Sparfels, noposplat], using higher-dimensional primitives [ndim], spatiotemporal models [4d], in addition to several methods to improve density control [3dgsmcmc, steepgs, pixelgs], anti-alising [mipsplat, anal, AA-2DGS], model compactness [compact, contextgs] and training speed [newton, grendel, speedysplat]. The 2DGS representation [2DGS] leverages planar 2D primitives instead of volumetric ones (3DGS), and performs precise 2D kernel evaluation in object space as opposed to approximative ones in screen space (3DGS), thus leading to superior geometric modelling and multi-view consistency.
2.2 Specular Reflection Modeling for Reflective Scenes
@@
Both vanilla NeRFs and 3DGS assume low-frequency view dependency. Hence, they can struggle with highly reflective scenes. One strategy to improve in this department is using shading functions that are reflection direction aware [ref_nerf, Ref-NeuS, jiang2024rethinking, nerv, nero]. For instance, Ref-NeRF [ref_nerf] extends NeRFs with a new parameterization for view-dependent radiance and incorporates normal vector regularization. The shading function can be more physically grounded, and this enables additional application such as relighting and material editing. In this regard, other NeRF and GS based methods proposed to model light interaction using the explicit rendering equation with BRDF functions. ENVIDR [envidr] Uses environment maps to capture spatially-varying reflections in neural rendering. The next wave of work tackled shading for Gaussian Splatting (e.g. [3iGS, GS-ROR, GS-IR, Gshader, 3DGS-DR, R3DG, RefGaussian]). GShader [Gshader] applies a simplified shading function on each fragment. 3DGS-DR [3DGS-DR] introduces deferred shading at pixel level, and stabilizes the optimization by smoothing out normal gradients.
Building on the 2DGS representation, our baseline Ref-Gaussian [RefGaussian] decomposes the scene into geometry, material and lighting through the split-sum approximation of the rendering equation while incorporating an indirect light attribute, enabling inter-reflections while being fast to render. It achieves the state-of-the-art performance on the standard reflective scene novel view synthesis benchmarks. We propose to enhance its physical material and normal representations through the use of per-primitive textures.
2.3 Texture Attributes in Gaussian Splatting
@@
Several works [gstex, BBSplat, texturedgaussiansenhanced3d, SuperGaussians, GaussianBillboards, HDGS] recently introduced the texture attribute representation for Gaussian Splatting as well. They use it to model a view independent component of the color. Differently, we explore this representation for rendering material properties and normals within 2DGS enabled physically based rendering to reconstruct challenging highly reflective scenes. We also propose leveraging texture atlases to enable hardware acceleration at test-time rendering. Closest to our context, concurrent work [GTAvatar] manages to encode spatially varying primitive material attributes in a single compact texture map thanks to their pre-fitted template mesh.
3 Method
@@
In this section, we present our approach to enhance 2D Gaussian Splatting (2DGS) [2DGS] for representing highly reflective 3D scenes. We first provide background on 2DGS, then introduce our per-primitive texture mapping method that enables high-frequency detail on flat Gaussian primitives. Finally, we describe our hardware-accelerated implementation using texture atlases and the physically-based rendering model we use for reflective scenes.
3.1 Background: 2D Gaussian Splatting
@@
We build upon 2D Gaussian Splatting (2DGS) [2DGS], which represents a scene using oriented planar disks offering improved multi-view consistency compared to 3D Gaussian Splatting [3dgs]. Each primitive is characterized by its position , tangential vectors and , and scaling factors . The primitive’s normal is defined by the cross product .
@@
For a point in the local coordinate space corresponding to a given pixel (i.e. ray-splat intersection), the 2D Gaussian value is evaluated as:
@@
3.2 Per-Primitive Texture Mapping
@@
3.3 Texture Mapping Implementation
@@
3.4 Hardware Acceleration via Texture Atlases
To efficiently render optimized scenes at test time with per-primitive textures, we leverage hardware-accelerated texture filtering by packing individual primitive textures into texture atlases. This approach is inspired by methods like Ptex [Burley2008Ptex] and seamless texture atlases [purnomo2004seamless], adapted to the specific needs of Gaussian splatting. We provide details about Texture Atlas construction and texture sampling in the Supplementary Material (Section 1).
3.5 Physically-Based Deferred Rendering
Following the deferred rendering approach used in 3DGS-DR [3DGS-DR] and Ref-Gaussian [RefGaussian], we first splat material attributes to screen-space buffers, then apply physically-based shading in a separate pass.
3.5.1 Material Properties
Each 2D Gaussian is associated with texture maps for the following material properties: Albedo , Metallic , Roughness , Tangent normal representing normal perturbations in the tangent space. For memory efficiency, we encode tangent normals using only two components and reconstruct the third component at runtime using .
3.5.2 Normal Mapping
3.5.3 Attribute Splatting
where represents the combined screen-space buffers:
This deferred approach treats alpha-blending as a smoothing filter, stabilizing the optimization of features sampled from textures and producing more cohesive rendering results compared to shading directly on the Gaussians [RefGaussian, 3DGS-DR].
3.5.4 Physically-Based Shading
The diffuse term writes:
The first term, , depends solely on the view angle and roughness, which is precomputed and stored in a 2D lookup texture. is the direct environment lighting queried from a learnable environment map in the reflection direction , using roughness for mipmap selection. is the blended indirect lighting component. We follow the baseline method [RefGaussian] in modeling inter-reflections by approximating visibility with ray tracing an extracted mesh and encoding indirect lighting with spherical harmonics. We follow prior work [RefGaussian] and use [NVDiffRec] for the PBR shading.
3.6 Training
where is the RGB reconstruction loss with balancing weight . The normal consistency loss encourages alignment of the Gaussians with the surface by minimizing the cosine difference between the rendered normal and the surface normal derived from rendered depth. We use a single NVIDIA RTX A6000 GPU in our experiments.
4 Experiments
| Shiny Blender [ref_nerf] | Glossy Synthetic [nero] | Real [ref_nerf] | ||||||||||||||||
@@ 4 Experiments
| LPIPS | Ours | 0.074 | 0.028 | 0.075 | 0.042 | 0.004 | 0.063 | 0.038 | 0.064 | 0.034 | 0.045 | 0.040 | 0.055 | 0.041 | 0.055 | 0.283 | 0.257 | 0.27 |
Following the baseline method [RefGaussian], we evaluate our work quantitatively and qualitatively under standard multi-view reconstruction benchmarks of challenging reflective scenes. We use the datasets: Shiny Blender [ref_nerf] and Glossy Synthetic [nero] for novel view synthesis of reflective objects, and dataset Ref-Real [ref_nerf] to account for real world open reflective scenes. We also provide results on the Synthetic NeRF [nerf] dataset to showcase our method under non-reflective scenes and illustrate its practicality in the supplementary material (Section 2). We compare to state-of-the-art methods in the reflective scene setting, including the baseline Ref-Gaussian [RefGaussian], other reflective Gaussian splatting based methods GShader [Gshader] and 3DGS-DR [3DGS-DR], 3DGS [3dgs] and 2DGS [2DGS] for reference, and seminal NeRF based approaches such as Ref-NeRF [ref_nerf] and ENVIDR [envidr]. We provide additional results and ablation studies in the supplementary material.
4.1 Implementation Details
We follow a two-stage optimization for stability. We first train for half the total number of iterations using per-splat single attribute optimization. During the second stage, we start optimizing the material and normal textures initialized from the corresponding attributes from the first stage, and we freeze the positions of the primitives. In all our evaluations, we use texture resolution of . We use the same hyperparameters and training strategies defined by our baseline Ref-Gaussian [RefGaussian]. We also follow the latter in replacing the integrated diffuse lighting by a spherical harmonics view dependent color for better fitting in the reflective setting. We implement efficient CUDA kernels on top of 2DGS and Ref-Gaussian for forward and backward operations involving material textures and normal mapping, as well as the texture atlases construction, packing and hardware bilinear filtering at test time.
4.2 Novel View Synthesis
Table 1 shows numerical results in the standard reflective benchmark. We report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [wang2004image], and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable] as metrics. Our method performs favorably compared to other methods, including baseline Ref-Gaussian, which is also the state-of-the-art method currently under this benchmark to the best of our knowledge. We provide qualitative results to accompany this table, for real scenes in Figure 3 and synthetic ones in Figure 2. Our method displays superior ability in capturing reflections on the surface with more fidelity, while Ref-Gaussian and 3DGS-DR suffer from distorted or missing reflections. We also show in the supplementary material numerical (Tab. 1 in Supp. Mat.) and qualitative (Fig. 1 in Supp. Mat.) comparisons under non-reflective data on Nerf Synthetic Scenes [nerf]. In that generic setting, we perform competitively with the state-of-the-art. The improvement brought by our method over the baseline Ref-Gaussian particularly, under both reflective and non-reflective scenes, is a testimony of the efficacy and versatility of our representation.
4.3 Scene Decomposition
Figure 4 shows a comparison of the decomposition with respect to our baseline. Notice that our material properties are sharper and display less noise. Our normals replicate the smooth sphere shape more faithfully. We provide further comparisons in supplementary material including environment map estimation and material decomposition (Figures 3 & 2 in Supp. Mat.) showcasing the superiority of our results.
We further evaluate our normal estimation through the benchmark of the Shiny Blender dataset [ref_nerf]. Table 2 reports the mean angular error of normal maps, where we outperform our baseline. This result validates our tangential normal representation and the use of normal mapping. Figure 5 shows qualitative comparisons of normals, where we recover superior geometry compared to other methods.